The length of the garden is 5 cm longer than its width and the area is 36 square meter. How long is the garden?
i need the also the solution and also the first question that i sent a while ago. Thank you :)

Answer by harpazo(655)

length = x + 5

width = x

Area = 36

Use A = length times width

36 = (x+5)(x)

36 = x^2 + 5x

x^2 + 5x - 36 = 0

Factor left side.

(x + 9)(x - 4) = 0

Set each factor to zero and solve for x.

REJECT the negative answer since we are looking for length.
x + 9 = 0
x = -9...REJECT this x value.
x - 4 = 0
x = 4...ACCEPT this value of x.

The length is x + 5 or 4 + 5 = 9cm.
The garden is 9cm long.